OpenStudy (anonymous):

1/5(2x-15)+3/5x=4 1/4

OpenStudy (anonymous):

What is x?

OpenStudy (paxpolaris):

\[\frac15\left( 2x-15 \right)+\frac35x=4\frac14\] is this correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So like what is x? I don't understand...

OpenStudy (paxpolaris):

you could multiply everything by 20 to get rid of the fractions:

OpenStudy (paxpolaris):

\[\frac 15 \times 20 = 4\] \[\frac 35 \times 20 = 4\] \[4\frac 14 \times 20 = {17\over4}\times20=17 \times 5 = 85\]

OpenStudy (paxpolaris):

**\[\frac 35 \times 20 = 12\]

OpenStudy (paxpolaris):

\[4(2x-15)+12x=85 \]

OpenStudy (paxpolaris):

\[8x-60+12x=85\]

OpenStudy (paxpolaris):

\[20x=85+60\]\[\implies 20x=145\]\[\implies x= {145 \over 20}={29 \over 4}=\color{green}{7\frac14}\]